# Holistic Reasoning with Long-Context LMs: A Benchmark for Database Operations on Massive Textual Data

### v1

@@ Abstract
The rapid increase in textual information means we need more efficient methods to sift through, organize, and understand it all. While retrieval-augmented generation (RAG) models excel in accessing information from large document collections, they struggle with complex tasks that require aggregation and reasoning over information spanning across multiple documents–what we call holistic reasoning. Long-context language models (LCLMs) have great potential for managing large-scale documents, but their holistic reasoning capabilities remain unclear. In this work, we introduce HoloBench, a novel framework that brings database reasoning operations into text-based contexts, making it easier to systematically evaluate how LCLMs handle holistic reasoning across large documents. Our approach adjusts key factors such as context length, information density, distribution of information, and query complexity to evaluate LCLMs comprehensively.
@@ 1 Introduction
The explosive growth of textual data has highlighted the need for efficient ways to process, organize, and comprehend large document collections. While retrieval-augmented generation (RAG) models have made accessing information from these vast sources easier (Lewis et al., 2020; Maekawa et al., 2024), these models are limited by their reliance on localized context retrieval. This makes them less effective for complex tasks that require holistic reasoning—the ability to synthesize, aggregate, and reason over information that spans across multiple documents. For instance, when asked, ”Which company employed the most people?”, traditional retrieval models might provide individual statistics for each company but fail to offer a comprehensive comparison across all relevant entities. This reveals a fundamental limitation: RAG systems excel at retrieving specific information but struggle to aggregate and reason over broader cross-document contexts. This gap highlights the need for more advanced models capable of holistic reasoning, particularly in tasks requiring multi-document synthesis, comparative analysis, and contextual integration across extensive datasets.
@@ 3.3 Implementation Details
Table 2 shows example queries for this categorization. Note that a single query can be associated with multiple types. These query types are automatically decided based on the type of database operator(s) in the SQL query. We also define three levels of query difficulty: Easy, Medium, and Hard. These are automatically decided based on the number of SQL components. Table 3 shows example queries by difficulty level. We provide more detailed definitions in Appendix A.2.1 and A.2.2. To observe the impact of information positioning, we consider three clustered positioning strategies –Beginning, Middle and End– and two distributed positioning strategies –Uniform and Bimodal (concentrated at both ends). These are illustrated in Figure 2.
@@ 4.1 Setup
We use nine recent LCLMs that support $128k$ or longer context length: Llama 3.1 8B/70B/405B-instruct (Dubey et al., 2024), GPT-4o-mini/GPT-4o (OpenAI, 2024a), o1-mini (OpenAI, 2024b), Claude 3.5 Sonnet (Anthropic, 2024), and Gemini 1.5 Flash/Pro (Reid et al., 2024).111gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, o1-mini-2024-09-12, claude-3-5-sonnet@20240620, gemini-1.5-flash-001, and gemini-1.5-pro-001 We used the same prompt across all the models shown in Appendix.
@@
Table 5: Model performances for different context lengths under both constant information amount and constant information density settings. Avg. refers to the average across context sizes, and wAvg. denotes the weighted average, with weights either increasing (inc) or decreasing (dec) linearly with context length. The best score of each column is underlined. Generally, performance declines as context length increases. The drop is wider for information density than information amount.
@@ 4.2.2 Effect of Information Position
Llama-3.1-405b maintains stable performance across various information positions, showing resilience whether the information is distributed uniformly or clustered. In contrast, GPT-4o performs better when relevant information is grouped at the beginning or middle, but its accuracy declines when the information is scattered or placed toward the end, indicating a preference for more structured information placement.
No More Lost-in-the-Middle
An interesting finding from our experiments is that both models performed best when the information was placed in the middle of the context. One seminal work reported that LCLMs often missed crucial information located in the middle (Liu et al., 2024). However, it appears that recent LLMs have been optimized to counter this weakness, enabling them to process mid-context information more effectively, which likely explains their improved performance in this scenario.
@@ 4.2.3 Effect of Query Type and Difficulty
Interestingly, GPT-4o performs on par with other query types for Aggregation queries at 4k context length. However, as the context length increases, there is a notable drop in performance, indicating the model struggles to aggregate relevant information effectively in longer contexts. To investigate this behavior, we select a random aggregation query: “Find the destination airport and its country that receives flights from the highest number of distinct source airports, and include that number in the result.” and inspect the model predictions. We find that GPT-4o correctly lists all relevant information and produces the correct answer when the context length is 4k. At 8k context length, it can find relevant information but begins to struggle with correctly retrieving and aggregating the information. Specifically, it returns the correct airports but miscounts the number of airports. At 64k context length, it skips the detailed analysis of the content and instead jumps straight to answering the question, leading to incorrect responses. We provide model outputs are in Figures 14, 15, and 16 in Appendix. This highlights the limitation of LCLMs’ step-by-step reasoning abilities as the context size increases.
@@
Query complexity often affects performance more than the total volume of information. Figure 5 summarizes the performance of Llama-3.1-405b and GPT-4o across different query difficulties. We observe that the relative order of difficulty—easy, medium, and hard—remains consistent across different context lengths, highlighting the alignment between the inherent complexity of the queries and the models’ capacity for holistic reasoning over extended contexts. However, hard queries consistently show the lowest performance across all context lengths, with a steep decline as context length extends, indicating the increasing challenge of reasoning over longer contexts with complex queries. Upon close inspection of failure cases, we observe that even though models can think step-by-step, they struggle to retrieve information for each condition for multi-conditional hard queries. Additionally, we find that when the volume of relevant information is too large, LCLMs complain about the input format being inappropriate and suggest using programmatic parsing. This points to an important research gap in the reasoning capability of LCLMs that demands further research. We observe similar trends for other models and provide further details in the Appendices A.2.1 and A.2.2.
@@ 4.3 Discussion
A promising future direction would be to develop a dynamic mechanism for determining the optimal amount of information to retrieve based on the query and context length, particularly when working with weaker models for holistic reasoning.
@@ 5 Conclusion
We have made several efforts to ensure the reproducibility of our work. First, the design principles, methodology, and evaluation setup for HoloBench are detailed in Sections 4 and 5. The SQL queries, natural language counterparts, and data preprocessing steps are provided in the Appendix (Sections A.1 and A.2.3). Moreover, the benchmark creation process, which includes database construction and verbalization steps, is clearly outlined in Section 4.2, and additional dataset details are available in the supplementary material.
@@ Appendix A Appendix
A.1 Dataset Details
A.1.1 Database Preprocess
@@
A.1.2 Examples of Dummy Documents
@@
A.2 Query
A.2.1 Query Type
@@ A.2.2 Query Difficulty
A.2.3 Query Examples
@@
You’ll be given a set of sentences to read through carefully. Once you’ve reviewed them, I’ll ask you a question related to the information in those sentences. Your job is to think critically about the details, analyze the sentences in relation to the question, and then provide your answer. If the information clearly supports a partial answer, provide that. However, if the evidence is unclear or insufficient, it is okay to respond with ”No answer.” **Input:** - **Sentences:** ‘‘‘
@@
- **Reasoning:** - [Describe how you thought through the sentences and how they helped you reach your conclusion. If the evidence is unclear or insufficient to provide a reliable answer, explain why. Your reasoning should not exceed 10,000 words.] - **Answer:** [Provide an answer only if it is clearly supported by the information in the sentences. If the evidence is unclear or insufficient, respond with ”No answer.”]
@@
You’ll be given a set of sentences to review carefully. Once you’ve reviewed them, I’ll ask you a question related to the information in those sentences. Your task is to provide a direct, clear answer based only on what is clearly supported by the information. If the evidence does not fully support the answer, provide a partial answer. If the information is unclear or insufficient, respond with ”No answer.” Do not provide explanations or reasoning. **Input:** - **Sentences:** ‘‘‘
@@
- **Answer:** [Provide only the direct final answer. If the information supports only part of the answer, provide a partial answer. If the evidence is unclear or insufficient, respond with ”No answer.” Do not include reasoning.]
@@
You will be given a question along with a response generated by an assistant and the corresponding ground truth data. Your task is to assess the response based on its accuracy and completeness in comparison to the ground truth. For each entry in the ground truth, determine whether the information provided by the assistant is an “Exact Match,” a “Partial Match,” or a “No Match.” #### **Evaluation Criteria:** - **Exact Match**: The assistant’s response precisely matches the ground truth in both content and detail. - **Partial Match**: The assistant’s response includes some correct information but is either incomplete, incorrectly ordered, or contains inaccuracies. - **No Match**: The assistant’s response does not accurately reflect the ground truth or is missing entirely. #### **Special Cases:** **Ground Truth is None**: - If the ground truth is None (represented as an empty list []): - **Exact Match**: If the assistant’s response indicates that there is no information or content. - **No Match**: If the assistant’s response provides any information when the ground truth is None. #### **Output Format:** - The output should be a list of objects where each object contains: - An “id” that matches the id of the corresponding ground truth entry. - A “label” indicating whether the assistant’s response is an “Exact Match”, “Partial Match”, or “No Match”. - The number of output objects should match the number of entries in the ground truth. ### **Examples:** {Several Examples} ====== Your task starts here ====== **Question:** ‘‘‘
@@
A.3 Prompt Templates
First, we provide a prompt used for our main experiments in Figure 8. Since o1-mini is a reasoning model, we remove the reasoning output part from the prompt for this model. Next, we show a prompt used for the “without CoT” setting in Figure 9. Also, Figure 10 shows an evaluation prompt.
A.4 Additional Experimental Results
@@
A.4.1 Information Position
To generalize our insights described in Section 4.2.2, we additional conducted experiments for smaller models, GPT-4o-mini, Gemini-1.5 Flash, and Llama-3.1-8b, with all information positions and report results in Figure 11. We observe that smaller models also have their preferences of information positions. For example, GPT-4o-mini achieves better performance on “middle” and “end” positions than other positions. For Gemini-1.5 Flash, “beginning” and “middle” positions are better than other positions. Interestingly, Llama-3.1-8b achieves better performance not only “beginning” and “middle” but also “uniform” positions. Since this preference on “uniform” is shared with its larger model Llama-3.1-405b, their training process enhance models’ information retrieval ability even if information pieces are distributed over documents.

### v2

@@ Abstract
The rapid increase in textual information means we need more efficient methods to sift through, organize, and understand it all. While retrieval-augmented generation (RAG) models excel in accessing information from large document collections, they struggle with complex tasks that require aggregation and reasoning over information spanning across multiple documents–what we call holistic reasoning. Long-context language models (LCLMs) have great potential for managing large-scale documents, but their holistic reasoning capabilities remain unclear. In this work, we introduce HoloBench, a novel framework a novel framework that brings database reasoning operations into text-based contexts, making it easier to systematically evaluate how LCLMs handle holistic reasoning across large documents. Our approach adjusts key factors such as context length, information density, distribution of information, and query complexity to evaluate LCLMs comprehensively.
@@ 1 Introduction
The explosive growth of textual data has highlighted the need for efficient ways to process, organize, and comprehend large document collections. While retrieval-augmented generation (RAG) models have made accessing information from these vast sources easier (Lewis et al., 2020; Maekawa et al., 2024), these models are limited by their reliance on localized context retrieval. This makes them less effective for complex tasks that require holistic reasoning—the ability to synthesize, aggregate, and reason over information that spans across multiple documents. For instance, when asked, "Which company employed the most people?", traditional retrieval models might provide individual statistics for each company but fail to offer a comprehensive comparison across all relevant entities. This reveals a fundamental limitation: RAG systems excel at retrieving specific information but struggle to aggregate and reason over broader cross-document contexts. This gap highlights the need for more advanced models capable of holistic reasoning, particularly in tasks requiring multi-document synthesis, comparative analysis, and contextual integration across extensive datasets.
@@ 3.3 Implementation Details
Table 2 shows example queries for this categorization. Note that a single query can be associated with multiple types. These query types are automatically decided based on the type of database operator(s) in the SQL query. We also define three levels of query difficulty: Easy, Medium, and Hard. These are automatically decided based on the number of SQL components. Table 3 shows example queries by difficulty level. We provide more detailed definitions in Appendix A.3.1 and A.3.2. To observe the impact of information positioning, we consider three clustered positioning strategies –Beginning, Middle and End– and two distributed positioning strategies –Uniform and Bimodal (concentrated at both ends). These are illustrated in Figure 2.
@@ 4.1 Setup
We use nine LCLMs that support $128k$ or longer context length: Llama 3.1 8B/70B/405B-instruct (Dubey et al., 2024), GPT-4o-mini/GPT-4o (OpenAI, 2024a), o1-mini (OpenAI, 2024b), Claude 3.5 Sonnet (Anthropic, 2024), Gemini 1.5 Flash/Pro (Reid et al., 2024) and five more recent LCLMs: o3-mini (OpenAI, 2025), Gemini 2.0 Flash/Flash-Thinking, and DeepSeek-V3/R1 (DeepSeek-AI, 2024; 2025). We used the same prompt across all the models shown in Appendix A.4.
@@
Gemini-2.0 Flash $66.2$ $62.3$ $59.1$ $57.3$ $53.0$ $59.6$ $57.5$ $61.7$ $66.2$ $60.1$ $57.0$ $52.5$ $48.2$ $56.8$ $53.9$ $59.7$
Gemini-2.0 Flash Thinking $69.5$ $54.2$ $59.5$ $59.6$ $43.0$ $57.2$ $54.0$ $60.3$ $69.5$ $55.3$ $51.9$ $43.2$ $30.3$ $50.1$ $44.0$ $56.1$
o3-mini $68.6$ $62.6$ $51.6$ $46.8$ $42.7$ $54.5$ $50.0$ $59.0$ $68.6$ $58.0$ $47.2$ $33.6$ $22.4$ $46.0$ $38.2$ $53.8$
DeepSeek-V3 $59.5$ $51.7$ $50.8$ $47.3$ $53.6$ $52.6$ $51.5$ $53.7$ $59.5$ $56.9$ $52.5$ $45.8$ $36.0$ $50.1$ $46.2$ $54.0$
DeepSeek-R1 $51.1$ $41.8$ $34.1$ $24.1$ $36.4$ $37.5$ $34.4$ $40.7$ $51.1$ $60.4$ $31.9$ $24.6$ $11.3$ $35.9$ $28.2$ $43.6$
Table 5: Model performances for different context lengths under both constant information amount and constant information density settings. Avg. refers to the average across context sizes, and wAvg. denotes the weighted average, with weights either increasing (inc) or pydecreasing (dec) linearly with context length. The best score of each column is underlined. Generally, performance declines as context length increases. The drop is wider for information density than information amount.
@@ 4.2.1 Effect of context length and information amount
Reasoning-focused LCLMs struggle with extended contexts
Reasoning-focused models—o1-mini, o3-mini, Gemini-2.0 Flash Thinking, and DeepSeek-R1—show strong performance on shorter contexts but experience significant performance degradation as context length increases. This suggests that current training approaches for reasoning-focused models may not adequately address the complexities of processing longer contexts. While general-purpose models like Gemini-2.0 Flash maintain consistent performance across context lengths, the challenges faced by reasoning-focused models highlight an important area for future research: developing training strategies that can maintain robust reasoning capabilities even with extended contexts.
@@ 4.2.2 Effect of Information Position
Llama-3.1-405b maintains stable performance across various information positions, showing resilience whether the information is distributed uniformly or clustered. In contrast, GPT-4o performs better when relevant information is placed at the end or bimodal, but its accuracy declines when the information is grouped at the beginning and middle, suggesting that it tends to favor details presented at the end.
@@ 4.2.3 Effect of Query Type and Difficulty
Interestingly, GPT-4o performs on par with other query types for Aggregation queries at 4k context length. However, as the context length increases, there is a notable drop in performance, indicating the model struggles to aggregate relevant information effectively in longer contexts. To investigate this behavior, we select a random aggregation query: “Find the destination airport and its country that receives flights from the highest number of distinct source airports, and include that number in the result." and inspect the model predictions. We find that GPT-4o correctly lists all relevant information and produces the correct answer when the context length is 4k. At 8k context length, it can find relevant information but begins to struggle with correctly retrieving and aggregating the information. Specifically, it returns the correct airports but miscounts the number of airports. At 64k context length, it skips the detailed analysis of the content and instead jumps straight to answering the question, leading to incorrect responses. We provide model outputs are in Figures 14, 15, and 16 in Appendix. This highlights the limitation of LCLMs’ step-by-step reasoning abilities as the context size increases.
@@
Query complexity often affects performance more than the total volume of information. Figure 5 summarizes the performance of Llama-3.1-405b and GPT-4o across different query difficulties. We observe that the relative order of difficulty—easy, medium, and hard—remains consistent across different context lengths, highlighting the alignment between the inherent complexity of the queries and the models’ capacity for holistic reasoning over extended contexts. However, hard queries consistently show the lowest performance across all context lengths, with a steep decline as context length extends, indicating the increasing challenge of reasoning over longer contexts with complex queries. Upon close inspection of failure cases, we observe that even though models can think step-by-step, they struggle to retrieve information for each condition for multi-conditional hard queries. Additionally, we find that when the volume of relevant information is too large, LCLMs complain about the input format being inappropriate and suggest using programmatic parsing. This points to an important research gap in the reasoning capability of LCLMs that demands further research. We observe similar trends for other models and provide further details in the Appendix A.5.
@@ 4.3 Discussion
Further, detailed comparisons between RAG and LCLMs are provided in Appendix A.5.5. A promising future direction would be to develop a dynamic mechanism for determining the optimal amount of information to retrieve based on the query and context length, particularly when working with weaker models for holistic reasoning.
@@ 5 Conclusion
We have made several efforts to ensure the reproducibility of our work. First, the design principles, methodology, and evaluation setup for HoloBench are detailed in Sections 4 and 5. The SQL queries, natural language counterparts, and data preprocessing steps are provided in the Appendix (Sections A.2 and A.3.3). Moreover, the benchmark creation process, which includes database construction and verbalization steps, is clearly outlined in Section 4.2, and additional dataset details are available in the supplementary material.
@@ Appendix A Appendix
A.1 Model Details
We show the model details in Table 6.
Model Size Context HuggingFace / API
GPT-4o (OpenAI, 2024a) - 128k gpt-4o-2024-08-06
GPT-4o-mini (OpenAI, 2024a) - 128k gpt-4o-mini-2024-07-18
o3-mini (OpenAI, 2025) - 200k o3-mini-2025-01-31
o1-mini (OpenAI, 2024b) - 128k o1-mini-2024-09-12
Claude-3.5-Sonnet (Anthropic, 2024) - 200k claude-3-5-sonnet@20240620
Gemini-1.5-Pro (Reid et al., 2024) - 2M gemini-1.5-pro-001
Gemini-1.5-Flash (Reid et al., 2024) - 1M gemini-1.5-flash-001
Gemini-2.0-Flash (Reid et al., 2024) - 1M gemini-2.0-flash-001
Gemini-2.0-Flash-Thinking (Reid et al., 2024) - 1M gemini-2.0-flash-thinking-exp-01-21
Llama-3.1-405b (Dubey et al., 2024) 405B 128k meta-llama/Llama-3.1-8B-Instruct
Llama-3.1-70b (Dubey et al., 2024) 70B 128k meta-llama/Llama-3.1-70B-Instruct
Llama-3.1-8b (Dubey et al., 2024) 8B 128k meta-llama/Llama-3.1-405B-Instruct
DeepSeek-V3 (DeepSeek-AI, 2024) 671B 128k deepseek-ai/DeepSeek-V3
DeepSeek-R1 (DeepSeek-AI, 2025) 671B 128k deepseek-ai/DeepSeek-R1
Table 6: Models used in experiments. Model sizes are not publicly disclosed (-).
A.2 Dataset Details
A.2.1 Database Preprocess
@@
A.2.2 Dataset Selection
To create HoloBench for evaluating LCLMs, we required large tables capable of generating long contexts. The following criteria were applied to ensure diversity in both domains and answers:
  1. 1.
    Databases must have more than two tables and at least 500 rows in total.
  2. 2.
    Each selected database must belong to a distinct domain.
Among the databases in the SPIDER dataset, only 15 meet the row count requirement:
[’soccer_1’, ’wta_1’, ’baseball_1’, ’sakila_1’, ’flight_4’, ’formula_1’, ’college_2’, ’bike_1’, ’store_1’, ’chinook_1’, ’world_1’, ’csu_1’, ’flight_2’, ’wine_1’, ’car_1’]
From these, we selected five databases spanning different domains while also diversifying the number of tables and rows. The selected databases are summarized in Table 7.
Base DB # Tables # Total Rows
Wine_1 3 555
Store_1 4 2,548
College_2 9 3,554
Flight_4 3 20,989
Soccer_1 5 185,608
Table 7: Summary of Selected Databases for HoloBench
This selection balances domain variety, table count, and row diversity to create a robust evaluation benchmark.
A.2.3 Examples of Dummy Documents
@@
A.3 Query
A.3.1 Query Type
@@ A.3.2 Query Difficulty
Since our questions are manually derived from SQL queries, more complex SQL queries with additional conditions result in longer and more intricate questions. To quantify this, we calculated the average word count per question across different levels of SQL query difficulty. The summarized results are presented in Table 8.
SQL Difficulty Average Question Length (words)
Easy 10.8
Medium 13.8
Hard 17.0
Table 8: Average Question Length by SQL Query Difficulty
A.3.3 Query Examples
@@
You’ll be given a set of sentences to read through carefully. Once you’ve reviewed them, I’ll ask you a question related to the information in those sentences. Your job is to think critically about the details, analyze the sentences in relation to the question, and then provide your answer. If the information clearly supports a partial answer, provide that. However, if the evidence is unclear or insufficient, it is okay to respond with "No answer." **Input:** - **Sentences:** ‘‘‘
@@
- **Reasoning:** - [Describe how you thought through the sentences and how they helped you reach your conclusion. If the evidence is unclear or insufficient to provide a reliable answer, explain why. Your reasoning should not exceed 10,000 words.] - **Answer:** [Provide an answer only if it is clearly supported by the information in the sentences. If the evidence is unclear or insufficient, respond with "No answer."]
@@
You’ll be given a set of sentences to review carefully. Once you’ve reviewed them, I’ll ask you a question related to the information in those sentences. Your task is to provide a direct, clear answer based only on what is clearly supported by the information. If the evidence does not fully support the answer, provide a partial answer. If the information is unclear or insufficient, respond with "No answer." Do not provide explanations or reasoning. **Input:** - **Sentences:** ‘‘‘
@@
- **Answer:** [Provide only the direct final answer. If the information supports only part of the answer, provide a partial answer. If the evidence is unclear or insufficient, respond with "No answer." Do not include reasoning.]
@@
You will be given a question along with a response generated by an assistant and the corresponding ground truth data. Your task is to assess the response based on its accuracy and completeness in comparison to the ground truth. For each entry in the ground truth, determine whether the information provided by the assistant is an “Exact Match," a “Partial Match," or a “No Match." #### **Evaluation Criteria:** - **Exact Match**: The assistant’s response precisely matches the ground truth in both content and detail. - **Partial Match**: The assistant’s response includes some correct information but is either incomplete, incorrectly ordered, or contains inaccuracies. - **No Match**: The assistant’s response does not accurately reflect the ground truth or is missing entirely. #### **Special Cases:** **Ground Truth is None**: - If the ground truth is None (represented as an empty list []): - **Exact Match**: If the assistant’s response indicates that there is no information or content. - **No Match**: If the assistant’s response provides any information when the ground truth is None. #### **Output Format:** - The output should be a list of objects where each object contains: - An “id” that matches the id of the corresponding ground truth entry. - A “label" indicating whether the assistant’s response is an “Exact Match", “Partial Match", or “No Match". - The number of output objects should match the number of entries in the ground truth. ### **Examples:** {Several Examples} ====== Your task starts here ====== **Question:** ‘‘‘
@@
A.4 Prompt Templates
First, we provide a prompt used for our main experiments in Figure 8. Since o1-mini is a reasoning model, we remove the reasoning output part from the prompt for this model. Next, we show a prompt used for the “without CoT" setting in Figure 9. Also, Figure 10 shows an evaluation prompt.
A.5 Additional Experimental Results
@@
A.5.1 Information Position
To generalize our insights described in Section 4.2.2, we additional conducted experiments for smaller models, GPT-4o-mini, Gemini-1.5 Flash, and Llama-3.1-8b, with all information positions and report results in Figure 11. We observe that smaller models also have their preferences of information positions. For example, GPT-4o-mini achieves better performance on “beginning" and “bimodal" positions than other positions. For Gemini-1.5 Flash, “end" and “bimodal" positions are better than other positions. Interestingly, Llama-3.1-8b achieves better performance not only “end" and “bimodal" but also “uniform" positions. Since this preference on “uniform" is shared with its larger model Llama-3.1-405b, their training process enhance models’ information retrieval ability even if information pieces are distributed over documents.